You pay 168 dollars to play a lottery. A bin contains 91 balls labeled from 1 to 91. You draw a ball and you receive cash in the amount of 3 times the number shown on the ball. Let X be the number shown on the ball.


(a) Identify the probability distribution for the X.


Distribution = Parameter =    


(b) Calculate the probability that you get more money back than you have spent on the ticket.


P(you win) =    


(c) Calculate the average number that you will draw. Calculate the standard deviation for this number.

μ =    

σ =    


(d) Calculate the average amount that you will gain (winnings minus what you paid) by purchasing a ticket. Calculate the standard deviation associated to this number.


Average Gain (Winnings - paid) =    

SD of Gain =    


1
Expert's answer
2022-04-20T14:32:54-0400

a) We can select each ball with probability of "1/91\approx0.011"

"P(X)=0.011~for~X~from~1~to~91"

This is a uniform distribution.


"b)~P(winning)=P(3X>168)=P(X>56)=\\
=(91-56)/91=5/13\approx0.385"


"c)~\mu=\frac{1}{91}\sum_{k=1}^{91}k=\frac{1}{91}\frac{91(91+1)}{2}=46\\
\sigma^2=\frac{1}{90}\sum_{k=1}^{91}(k-\mu)^2=\frac{1}{90}(\sum_{k=1}^{45}(46-k)^2+\\
+\sum_{k=47}^{91}(k-46)^2)=\frac{1}{90}(\sum_{k=1}^{45}k^2+\sum_{k=1}^{45}k^2)=\\
=\frac{1}{45}\sum_{k=1}^{45}k^2"

For the sum of squares we will use the following formula:

"\sum_{k=1}^nk^2=\frac{1}{6}n(n+1)(2n+1)"

"\sigma^2=\frac{1}{45\cdot6}45\cdot46\cdot91\approx698\\
\sigma=\sqrt{\sigma^2}\approx26.4"


d) Average gain:

"G=\frac{1}{91}\sum_{k=1}^{91}(3k-168)=\\
=3\mu-168=-30"

Variance of gain:

"SD^2=\frac{1}{90}\sum_{k=1}^{91}((3k-168)-(-30))^2=\\
=\frac{1}{90}\sum_{k=1}^{91}(3k-138)^2=\frac{1}{90}(9\sum_{k=1}^{91}k^2-\\
-2\cdot3\cdot138\sum_{k=1}^{91}k+91\cdot138^2)=\\
=\frac{1}{90}(9\cdot\frac{1}{6}\cdot91\cdot92\cdot183-138(6\cdot\frac{91\cdot92}{2}-\\
-91\cdot138))=6279"

Standard deviation of gain:

"SD=\sqrt{SD^2}\approx79.24"